Mike drives to a friend's house at 40 mi/h. He drives home the same route at 45 mi/h. The total trip took 4 hours. What is the total distance mike drives?
Answer by ewatrrr(24785)

Hi,
*Note: Let t be the time to drive to Friend's house. (4hr-t) would then be the time left to drive home.
.
distance same both ways r*t = d
40(mi/hr)*t = 45(mi/hr)*(4hr-t)
.
Simplify and Solve
85(mi/hr)*t= 180mi
t = 180/85 hr
.
d =40(mi/hr)*180/85 hr = 84.7mi